OpenStudy (anonymous):

What is the probability of flipping a quarter 3 times and getting all heads?

OpenStudy (anonymous):

(1/2)^3 = 1/8

OpenStudy (the_fizicx99):

1/2 * 1/2 * 1/2 = (1/2)^3 = 1/8 though the odds of 1/8 doesn't make sense to me, I think i did something wrong

OpenStudy (mathmale):

What is the probability of obtaining HEADS on a single flip? Does the outcome of one flip affect the outcome of the next one? When no outcome depends upon any other outcome, how do you find the probability that you'll obtain the same outcome twice in a row? Three times in a row?